### v1

# Candidate Monotonicity and Proportionality for Lotteries and Non-Resolute Rules

Abstract
We study the problem of designing multiwinner voting rules that are candidate monotone and proportional. We show that the set of committees satisfying the proportionality axiom of proportionality for solid coalitions is candidate monotone. We further show that Phragmén’s Ordered Rule can be turned into a candidate monotone probabilistic rule which randomizes over committees satisfying proportionality for solid coalitions.
1 Proportionality and Monotonicity
Proportional multiwinner voting rules are among the most well studied in computational social choice. Among them, the Single Transferable Vote (STV) is used in several countries, e.g., Australia, Malta, or Scotland, to elect governing bodies. While STV can be considered proportional (Dummett, 1984; Tideman, 1995) one of its flaws is its failure of candidate monotonicity: strengthening a winning candidate can make this candidate stop winning (Woodall, 1997; McCune and Graham-Squire, 2024). Interestingly enough, whether this flaw is a necessity is an open question. Namely, it is unknown whether there is any voting rule satisfying the proportionality axiom of proportionality for solid coalitions (PSC) together with candidate monotonicity (Woodall, 1997; Aziz and Lee, 2020) with most known proportional voting rules failing candidate monotonicity already for very small instances sizes (Aziz and Lee, 2020). Despite Woodall already writing that “a major unsolved problem is whether there exist rules that retain the important political features of STV and are also more monotonic” (Woodall, 1997) no such rule has been found or proven to be nonexistent.
Our Contribution.
In this paper, we tackle this open question from two slightly weaker angles. Firstly, we study non-resolute voting rules: rules that do not just output one committee, but a whole set of committees. For this setting, we show that the rule outputting all committees satisfying PSC is indeed candidate monotone. Secondly, we study probabilistic voting rules, i.e., rules selecting a lottery over committees. For this, we show that a variant of Phragmén’s Ordered Rule (Janson, 2016) can be adapted into a candidate monotone lottery over committees satisfying PSC.
Related Work.
Our work is situated in the area of proportional multiwinner voting with ranked preferences, in particular in the axiomatic study of multiwinner voting rules. This study dates back to the work of eminent philosopher Dummett (1984) and his development of PSC. Aziz and Lee (2020) were the first to give slightly improved results for monotonicity and proportionality, showing that their Expanding Approvals Rule (EAR) satisfies a weaker form of monotonicity together with a generalization of PSC in the setting of weak ranked preferences in which ties are allowed. The same generalization of PSC together with a form of monotonicity they call “indifference monotonicity” is also satisfied by the generalization of STV by Delemazure and Peters (2024). McCune and Graham-Squire (2024) experimentally studied monotonicity violations in a real-world dataset consisting of Scottish local elections and found that out of 1079 elections 62 possessed some kind of “monotonicity anomaly”. In particular, in $23$ elections STV failed candidate monotonicity as studied by us.
PSC was further studied by Brill and Peters (2023) who criticized it for being a relatively weak axiom and introduced stronger proportionality axioms still satisfied by EAR. Further, Aziz and Lee (2021) introduced PSC generalizations for the context of participatory budgeting (Peters et al., 2021) and in Aziz and Lee (2022) provided a characterization of PSC in the standard ordinal preference setting. Finally, recently Aziz et al. (2024a) studied committee monotonicity, i.e., the selection of a ranking of candidates, together with PSC and showed that a rule satisfying both PSC and committee monotonicity always exists.
In the setting of approval-based multiwinner voting, monotonicity is significantly easier to obtain (Sánchez-Fernández and Fisteus, 2019) with most rules being monotone in the sense that voters additionally approving an already winning candidate cannot hurt that candidate. However, it can for instance happen that adding voters only approving a single candidate can make this candidate not be chosen anymore (see Boehmer et al. (2024) for examples of this in the context of participatory budgeting).
Finally, our results are also situated in the area of probabilistic social choice. Here, we are closely related to recent work focusing on best of both world guarantees, i.e., the design of lotteries satisfying desirable properties both ex-ante and ex-post. In particular, recent work focusing on the design of randomized apportionment schemes (Gölz et al., 2024; Correa et al., 2024) and on achieving ex-post and ex-ante proportionality guarantees in approval-based multiwinner voting and participatory budgeting (Aziz et al., 2023, 2024b; Suzuki and Vollen, 2024) are close to our section on probabilistic voting rules.
2 Preliminaries
Notation.
For $t\in\mathbb{N}$ let $[t]\coloneqq\{1,\dots,t\}$. We are given a set $N=[n]$ of voters and a set $C=\{c_{1},\dots,c_{m}\}$ of candidates. Each voter $i\in N$ possesses a strict linear ranking $\succ_{i}\subseteq C\times C$ over all candidates. Further, we are given a committee size $k\leq m$. Together, the voters, candidates, preferences, and committee size form an instance $\mathcal{I}=(N,C(\succ_{i})_{i\in N},k)$. A committee is a subset $W\subseteq C$ of size $\lvert W\rvert=k$. Let $C_{k}=\{W\subseteq C\colon\lvert W\rvert=k\}$ be the set of all committees. A rule $r$ is a function mapping an instance $\mathcal{I}$ to a non-empty set of committees $r(\mathcal{I})\subseteq C_{k}$. If $\lvert r(\mathcal{I})\rvert=1$ for all instances $\mathcal{I}$ the rule is said to be resolute.
For probabilistic rules, we mostly follow the notation of Aziz et al. (2023). We say that a fractional committee is an assignment of probabilities $(p_{1},\dots,p_{m})\in[0,1]^{m}$ for each candidate with $\sum_{i\in m}p_{i}=k$ with a fractional committee rule mapping instances to fractional committees. We let $\Delta(C_{k})$ be the set of all probability distributions over $C_{k}$. A probabilistic rule $r$ then maps an instance $\mathcal{I}$ to a probability distribution $r(\mathcal{I})\in\Delta(C_{k})$. Each probabilistic rule also has a natural non-resolute counterpart $\operatorname{supp}(r(\mathcal{I}))$, i.e., the set of committees being selected with non-zero probability (or the support of $r(\mathcal{I})$). Further, each probabilistic rule $r$ implements a fractional committee voting rule, with the probability of each candidate being the probability of this candidate being on the committee selected by $r$. On the other hand, there can be several probabilistic voting rules implementing a fractional committee voting rule.
Proportionality.
Following the seminal work of Dummett (1984) we define the axiom of proportionality for solid coalitions (PSC). For this, we first need to define the eponymous solid coalitions.
Definition 1 (Solid Coalition).
A group $N^{\prime}\subseteq N$ of voters forms a solid coalition over a group $C^{\prime}\subseteq C$ of candidates if
$C^{\prime}\succ_{i}C\setminus C^{\prime}\text{ for all }i\in N^{\prime}.$
In essence, PSC now requires that every solid coalition is represented proportionally.
Definition 2 (Proportionality for Solid Coalitions (Dummett, 1984)).
A committee $W$ satisfies proportionality for solid coalitions (PSC) if there is no group $N^{\prime}\subseteq N$ of voters, $C^{\prime}\subseteq C$ of candidates and $\ell\in[k]$ such that
  • i.)
    $N^{\prime}$ forms a solid coalition over $C^{\prime}$
  • ii.)
    $\lvert N^{\prime}\rvert\geq\ell\frac{n}{k}$
  • iii.)
    $\lvert W\cap C^{\prime}\rvert<\min(\lvert C^{\prime}\rvert,\ell)$.
In other words, if $C^{\prime}$ forms a prefix of the preferences of all voters in $N^{\prime}$ then $W$ needs to contain a large enough “proportional share” of them to satisfy $N^{\prime}$.
There are several rules satisfying PSC, among them the famous single-transferable vote (STV) (Tideman, 1995) and the expanding approvals rule (EAR) (Aziz and Lee, 2020).
Note that due to our definition of rules the set of all committees satisfying PSC is also a (non-resolute) rule.
Example 1.
Consider the following instance with four voters, four candidates, and $k=2$:
  • $c_{1}\succ c_{2}\succ c_{3}\succ c_{4}$
  • $c_{2}\succ c_{1}\succ c_{3}\succ c_{4}$
  • $c_{3}\succ c_{4}\succ c_{1}\succ c_{2}$
  • $c_{4}\succ c_{2}\succ c_{1}\succ c_{3}$.
This instance contains one non-trivial solid coalition large enough to be able to demand a candidate, namely the first and second voter over $\{c_{1},c_{2}\}$ (and technically also over $\{c_{1},c_{2},c_{3}\}$). Thus, any committee containing at least one one $c_{1}$ or $c_{2}$ satisfies PSC. Now if the third voter would swap candidate $c_{4}$ up the third and fourth voter would form a solid coalition over $\{c_{4}\}$. Therefore, in this new instance the only two PSC committees are $\{c_{1},c_{4}\}$ and $\{c_{2},c_{4}\}$
Monotonicity.
For our definition of (candidate) monotonicity for non-resolute voting rules, we follow Elkind et al. (2017).
Definition 3 (Candidate Monotonicity (Elkind et al., 2017)).
A rule $r$ satisfies candidate monotonicity if for every instance $\mathcal{I}$, committee $W\in r(\mathcal{I})$ and $c\in W$ in any instance $\mathcal{I}^{\prime}$ obtained by shifting $c$ one rank up in the preference list of any voter, there is still some committee $W^{\prime}\in r(\mathcal{I}^{\prime})$ with $c\in W^{\prime}$.
That is, if the candidate $c$ is in some winning committee and is “made stronger” $c$ is still in some winning committee afterward. It is known that neither STV nor EAR are candidate monotone (Aziz and Lee, 2020). However, Phragmén’s ordered rule is indeed candidate monotone (Janson, 2016).
To generalize candidate monotonicity to probabilistic rules, we instead focus on the probabilities of individual candidates.
Definition 4 (Candidate Monotonicity for Probabilistic Rules).
A probabilistic rule $r$ satisfies candidate monotonicity if for every instance $\mathcal{I}$ and $c\in C$ in any instance $\mathcal{I}^{\prime}$ obtained by shifting $c$ one rank up in the preference list of any voter it holds that
$\Pr[c\in r(\mathcal{I})]\leq\Pr[c\in r(\mathcal{I}^{\prime})].$
In other words, shifting a candidate up in the preference lists of voters does not decrease their chances to be selected. This is indeed in some sense a stronger requirement than candidate monotonicity for non-resolute rules.
Observation 1.
If $r$ is a candidate monotone probabilistic rule, then $\operatorname{supp}(r)$ is a candidate monotone non-resolute rule.
Similarly, one can also define candidate monotonicity for fractional voting rules by looking at the individual probabilities.
3 Non-Resoluteness and Monotonicity
We start off our results by showing that the rule that selects all PSC committees is candidate monotone. For this we crucially make use of a characterization of PSC by Aziz and Lee (2022) using their minimal demand rules. These works as follows: We start off with an empty committee $W=\emptyset$. For each rank $r=1,\dots,m$ we check if there is a group of voters forming a solid coalition over a set of candidates of size $r$. If, additionally, this solid coalition witnesses a PSC violation, we add a candidate from $C^{\prime}$ to $W$. We repeat this until there is no further PSC violation, afterwards we increment $r$.
A committee now satisfies PSC if and only if it can be obtained via this process.
Theorem 2 (Characterization of PSC (Aziz and Lee, 2022)).
A committee $W$ satisfies PSC if and only if it can be obtained by a minimal demand rule.
Using this characterization, we can now show that the set of PSC committees is indeed candidate monotone.
Theorem 3.
The rule selecting all PSC committees is candidate monotone.
Proof.
Let $W$ be a committee satisfying PSC and $c\in W$. Assume that there is voter $i$ swapping $c$ up in their preference list. As $W$ satisfies PSC in the original preferences, it can be obtained via a minimal demand rule. If $W$ still satisfies PSC, even after the preference change, we are already done. Otherwise, there must exist a solid coalition $N^{\prime}$ and $C^{\prime}$ witnessing a PSC violation. We now follow the characterization of Aziz and Lee (2020) and iteratively check for each rank if there is a solid coalition witnessing a PSC violation. If there is one and the solid coalition does not include $i$ and $c$ this solid coalition already exists in the original instance. Thus, this solid coalition must have appeared in the original minimal demand rule, and we can include the same candidate as was included in $W$. If the solid coalition includes $i$ and $c$ however, we can include $c$ as it is part of the candidates witnessing the violation. As $W$ does not satisfy PSC, such a violation must appear eventually. The rest of the committee can then be filled up with an arbitrary execution of the minimal demand rule, therefore leading to a committee satisfying PSC and including $c$. ∎
4 A Probabilistic Version of Phragmén’s Ordered Rule
While the previous result gives us some hope for obtaining a proportional and candidate monotone rule, it is also quite weak. There can be a lot of committees satisfying PSC (Bardal et al., 2024) and our monotonicity notion only guarantees us that some committee includes the strengthened candidate. To remedy this, we study probabilistic voting rules. In particular, we use a variant of the Phragmén’s Ordered Rule (Janson, 2016) and turn it into a lottery over committees satisfying PSC.
Phragmén’s Ordered Rule.
Phragmén’s Ordered Rule was originally introduced by a Swedish royal commission (including Phragmén himself) and was used for the intra party seat distribution in the Swedish parliament, for more historic context see the work by Janson (2016).
Phragmén’s Ordered Rule works as a continuous process. Each voter starts to “eat” their top choice candidate at the same speed. Once a candidate is fully eaten, this candidate is elected onto the committee and the voters move on to their top choice “uneaten” candidate. This is done until $k$ candidates are elected. [^2] This is very similar to the famous probabilistic serial method of Bogomolnaia and Moulin (2001) and the veto by consumption rule of Ianovski and Kondratev (2021). For a connection between these three methods and the problem of designing a voting rule with low metric distortion see the note by Peters (2023).
Example 2.
Consider the following instance with four voters, four candidates, and $k=2$:
  • $c_{1}\succ c_{4}\succ c_{3}\succ c_{2}$
  • $c_{1}\succ c_{3}\succ c_{2}\succ c_{4}$
  • $c_{2}\succ c_{3}\succ c_{4}\succ c_{1}$
  • $c_{3}\succ c_{4}\succ c_{2}\succ c_{1}$
In this instance Phragmén’s Ordered Rule works as follows:
  1. 1.
    At time $t=0$, all voters start “eating” their top choice candidate.
  2. 2.
    At time $t=0.5$:
    • $c_{1}$ is fully eaten and elected to the committee.
    • $c_{2}$ and $c_{3}$ are “half eaten”.
    • The first two voters move to their next choices $c_{4}$ ad $c_{3}$.
    • The current committee is $\{c_{1}\}$.
  3. 3.
    At time $t=0.75$:
    • $c_{3}$ is now fully eaten by the second and fourth voter, $c_{3}$ is $75\%$ eaten and $c_{4}$ is $25\%$ eaten.
    • $c_{3}$ is elected to the committee and $\{c_{1},c_{3}\}$ is the outcome.
To see that Phragmén’s Ordered Rule does not satisfy PSC consider the following example.
Example 3.
Consider the following instance with $15$ voters, $8$ candidates and $k=3$:
  • 2×2$\times$2 ×
    $c_{1}\succ c_{2}\succ c_{3}\succ c_{4}\succ c_{5}\succ c_{6}\succ c_{7}$
  • 2×2$\times$2 ×
    $c_{2}\succ c_{3}\succ c_{4}\succ c_{5}\succ c_{1}\succ c_{6}\succ c_{7}$
  • 2×2$\times$2 ×
    $c_{3}\succ c_{4}\succ c_{5}\succ c_{1}\succ c_{2}\succ c_{6}\succ c_{7}$
  • 2×2$\times$2 ×
    $c_{4}\succ c_{5}\succ c_{1}\succ c_{2}\succ c_{3}\succ c_{6}\succ c_{7}$
  • 2×2$\times$2 ×
    $c_{5}\succ c_{1}\succ c_{2}\succ c_{3}\succ c_{4}\succ c_{6}\succ c_{7}$
  • 5×5$\times$5 ×
    $c_{6}\succ c_{7}\succ c_{1}\succ c_{2}\succ c_{3}\succ c_{4}\succ c_{5}$
Here, Phragmén’s Ordered Rule would first elect $c_{6}$. Then at the time $c_{7}$ is fully eaten, every candidate from $c_{1}$ to $c_{5}$ is $80\%$ eaten. Thus, Phragmén’s Ordered Rule first elects $c_{6}$ and $c_{7}$ and only one of $c_{1}$ to $c_{5}$ afterwards. However, the first ten voters form a solid coalition over $c_{1}$ to $c_{5}$ demanding at least two candidates while they only receive one. Therefore, Phragmén’s Ordered Rule fails PSC in this instance.
We now turn this deterministic rule first into a fractional one and then via dependent rounding into a probabilistic one. Specifically, we first define a version of Phragmén’s Ordered Rule giving us probabilities for the individual candidates to be elected and then show how to round these probabilities to guarantee that every committee in the support satisfies PSC.
We assume that each candidate has a weight of one. Now instead of stopping when $k$ candidates have been selected, we stop once the total eaten weight is $k$. Candidates can now be eaten fractionally, with $\alpha$ of a candidate being eaten corresponding to an $\alpha$ probability of this candidate being selected. We call this rule Phragmén’s fractional rule (PFR).
Example 4.
Consider the following instance with four voters, four candidates, and $k=2$:
  • $c_{1}\succ c_{2}\succ c_{3}\succ c_{4}$
  • $c_{1}\succ c_{4}\succ c_{3}\succ c_{2}$
  • $c_{1}\succ c_{3}\succ c_{2}\succ c_{4}$
  • $c_{3}\succ c_{4}\succ c_{2}\succ c_{1}$
After the total eaten weight is $\frac{4}{3}$, candidate $c_{1}$ is eaten fully, while candidate $c_{3}$ is eaten a third. Then the first three voters move on to their second choice. After the total eaten weight is $2$ candidates $c_{2}$ and $c_{4}$ both have a weight of $\frac{1}{6}$, while $c_{3}$ has $\frac{2}{3}$. Thus, candidate $c_{1}$ is selected with probability $1$ while $c_{2}$ and $c_{4}$ are selected with probability $\frac{1}{6}$ and $c_{3}$ is selected with probability $\frac{2}{3}$.
In the instance in Example 3, $c_{6}$ would get a probability of $1$ while $c_{1}$ to $c_{5}$ each get a probability of $\frac{2}{5}$.
While PFR gives us probabilities, it does not actually give us a distribution over committees yet. For this, we turn to the dependent rounding scheme of Gandhi et al. (2006).
Theorem 4 ((Gandhi et al., 2006)).
Let $G=(V,E)$ be a bipartite graph with weights $w_{e}\in[0,1]$ for each $e\in E$. Then there exists a lottery over subsets of edges $(\lambda_{i},E_{i})_{E_{i}\subseteq E}$ where the subset of edges $E_{i}\subseteq E$ is selected with probability $\lambda_{i}$ such that (i) each edge is selected with probability $w_{e}$ and (ii) in each $E_{i}$ a vertex $v$ has at least a degree of $\lfloor\sum_{e\in N(v)}w(e)\rfloor$ and at most $\lceil\sum_{e\in N(v)}w(e)\rceil$.
That is given a bipartite graph with probabilities on edges, we can round these edges such that the original probabilities are respected and the degrees in the selected graph are the rounded degrees in the original weighted graph.
Before we turn to the rounding scheme, we additionally introduce another property satisfied by PFR. Namely, it satisfies an ex-ante version of PSC, which requires that any group of at least $\frac{\ell n}{k}$ voters (even for non-integer $\ell$) gets a probability mass of at least $\ell$.
Definition 5.
A fractional committee rule $p$ satisfies ex-ante PSC if there is no group $N^{\prime}\subseteq N$ of voters, $C^{\prime}\subseteq C$ of candidates and $\ell\leq k$ such that
  • i.)
    $N^{\prime}$ forms a solid coalition over $C^{\prime}$
  • ii.)
    $\lvert N^{\prime}\rvert\geq\ell\frac{n}{k}$
  • iii.)
    $\sum_{c\in C^{\prime}}p_{c}<\min(\lvert C^{\prime}\rvert,\ell)$.
Now we turn to the main result of this section and show that PFR is indeed candidate monotone and satisfies this ex-ante PSC notion. Further, we use the dependent rounding scheme of Gandhi et al. (2006) to show that the probabilities given by PFR can be decomposed into a probabilistic rule only randomizing over committees satisfying PSC.
Theorem 5.
PFR is candidate monotone, satisfies ex-ante PSC, and can be decomposed into a distribution over PSC committees.
Proof.
The monotonicity of PFR follows from the proof of Janson (2016, Theorem 14.3), as he shows that at any time, after a candidate gets swapped up, this candidate is always “at least as much eaten” as before swapping this candidate up. In particular this holds for time step $k$ and thus the monotonicity follows.
For the ex-ante PSC guarantee, consider any group of size $\ell\frac{n}{k}$ forming a solid coalition over $C^{\prime}$. After $k$ time steps, the group has eaten exactly $\ell$ parts of candidates. As $C^{\prime}$ forms a prefix of the preferences of these voters, it must hold that either $C^{\prime}$ is eaten in total, thus giving a probability mass of $\lvert C^{\prime}\rvert$ or that at least $\ell$ probability mass is spent on candidates in $C^{\prime}$ as voters in $N^{\prime}$ would first start eating candidates outside of $C^{\prime}$ once everything in $C^{\prime}$ is eaten.
For the ex-post guarantee, we make use of the dependent rounding scheme of Theorem 4 while following the general structure of the minimal demand rules of Aziz and Lee (2022).
For a candidate $c\in C$ and voter $i\in N$ let $p_{c}(i)\in[0,1]$ be the amount of $c$ eaten by $i$ during the execution of PFR. Further, let $p_{c}$ be the probability of candidate $c$ being selected by PFR.
We start off with the bipartite graph $G=(\emptyset\cup C,E)$. As a first step, for each candidate $c\in C$ with $p_{c}=1$ we add a voter $v_{c}$ to $G$ with an edge to $c$ of weight $1$ and set $p_{c}(i)=0$ for all $i\in V$. Now, we go from rank $r=1$ to $m$. If there is a group $N^{\prime}\subseteq N$ of voters forming a solid coalition with $C^{\prime}$ of size $\lvert C^{\prime}\rvert=r$, let $\ell=\left\lfloor\sum_{i\in N^{\prime}}\sum_{c\in C^{\prime}}p_{c}(i)\right\rfloor$. If $\ell=0$ we continue. Otherwise, let $p^{\prime}_{c}(i)$ for $c\in C^{\prime}$ and $i\in N^{\prime}$ such that $p^{\prime}_{c}(i)\leq p_{c}(i)$ and $\sum_{c\in C^{\prime},i\in N^{\prime}}p^{\prime}_{c}(i)=\ell$. We add a group voter $i^{\prime}$ to the left side of the bipartite graph and connect it to each candidate in $c\in C^{\prime}$ with probability $\sum_{i\in N^{\prime}}p^{\prime}_{c}(i)$. Further, we deduct $p_{c}(i)$ by $p^{\prime}_{c}(i)$. If there is no such group, we increase $r$.
The dependent rounding scheme now selects committees according to $G$.
We need to show two things, (i) it respects the probabilities and (ii) the committees output by the rounding scheme satisfy PSC.
For the first point, we see that the probability spent on a single candidate $c$ is at most $\sum_{v\in V}p_{c}(v)$ by definition. Further, since for $r=m$ all voters form a solid coalition, it is also exactly $\sum_{v\in V}p_{c}(v)$ since $\sum_{c^{\prime}\in C,v\in V}p_{c}(v)$ is integral and only integral quantities get deducted. Thus, at the final step, there is still some integral probability mass left leading to all probabilities being fulfilled. Since each candidate has a probability mass of at most one, no candidate is elected more than once.
For (ii) let $N^{\prime}$ be a solid coalition over $C^{\prime}$ such that $\lvert N^{\prime}\rvert\geq\frac{\ell n}{k}$ for some $\ell\in[k]$. For $r=\lvert C^{\prime}\rvert$ we know that $C^{\prime}$ forms a prefix of the preferences of $N^{\prime}$. Any candidate from $C^{\prime}$ who has a probability of $1$ of getting selected is already included in the committee. Hence, the only remaining candidates have a probability of less than $1$. Let $C^{\prime\prime}$ be these candidates. By the way PFR runs, we know that the voters do not eat any candidate ranked worse. Hence, since the $C^{\prime}$ form a prefix for all voters in $N^{\prime}$ and since at most $\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$ budget could have been spent on candidates with probability $1$, there is still a budget of at least $\frac{\lvert N^{\prime}\rvert k}{n}-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert\geq\ell-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$ left at the start of the process. In iteration $r$ this will lead to the creation of vertices on the left side of the bipartition with weight at least $\lfloor\frac{\lvert N^{\prime}\rvert k}{n}-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert\rfloor\geq\ell-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$. Since each vertex has integral weight, this means that either every candidate in $C^{\prime\prime}$ or at least $\ell-\lvert C^{\prime}\setminus C^{\prime\prime}\rvert$ candidates are selected from $C^{\prime}$ via dependent rounding, leading to at least $\min(C^{\prime},\ell)$ candidates being selected, therefore showing that PSC is satisfied. ∎
Adding on to this, we can however also show that PFR cannot be used to obtain the stronger guarantee of rank-PJR (Brill and Peters, 2023).
Example 5.
Consider Example 5 of Brill and Peters (2023) with $4$ voters, $6$ candidates, and $k=2$:
  • $c_{1}\succ c_{2}\succ c_{3}\succ c_{4}\succ c_{5}\succ c_{6}$
  • $c_{5}\succ c_{2}\succ c_{3}\succ c_{4}\succ c_{6}\succ c_{1}$
  • $c_{4}\succ c_{3}\succ c_{2}\succ c_{6}\succ c_{1}\succ c_{5}$
  • $c_{6}\succ c_{3}\succ c_{2}\succ c_{5}\succ c_{1}\succ c_{4}$.
Here PFR would place a probability mass of $\frac{1}{2}$ on each of $c_{1},c_{4},c_{5},c_{6}$. This, however, is not compatible with rank-PJR, as it requires at least one of $c_{2}$ or $c_{3}$ to be selected.
4.1 Consequences for Party List Elections
A special case of multiwinner voting are so-called party list elections (Brill et al., 2024) in which a candidate can not only be elected once but multiple times. Of special interest is the apportionment problem. In it, voters give their vote for a single party, which then is aggregated into seats for the parties. That is, there are $m$ parties, with party $i$ receiving a $p_{i}$ share of the votes and $h$ seats to be distributed among the parties. Just as for multiwinner voting, there are also “proportionality” axioms for apportionment. Most prominently, the quota axiom states that party $i$ should receive either $\lfloor p_{i}h\rfloor$ or $\lceil p_{i}h\rceil$ seats, i.e., the number of seats for each party should be correctly rounded. Since allocation rules satisfying quota and other prominent desirable axioms do not exist (Balinski and Young, 2001), recent work has studied the probabilistic allocation of seats (Grimmet, 2004; Gölz et al., 2024; Correa et al., 2024): the vote shares of the parties get probabilistically rounded in such a way that each party receives their exact quota in expectation and always their rounded quota ex-post. Here, we shortly note that we can apply our result on PFR to show that if one additionally knew the complete rankings of each voter over all parties, one can get a randomized apportionment scheme satisfying the desirable quota properties in addition to PSC (when interpreting this apportionment instance as a multiwinner voting instance).
Theorem 6.
There is a randomized apportionment method satisfying ex-ante exact quota, ex-post quota and PSC.
Proof.
For this, we note that PFR when run on a party-list instance, leads to probabilities exactly equal to the vote shares (as the candidates of a party never run out). Thus, Theorem 5 immediately gives us that ex-ante exact quota and PSC are satisfied. Further, due to the dependent rounding works, we also get that no party would be either over or underrepresented thus also giving us ex-post quota. ∎
[…]
Figure 2: On the left, the number of instances with a given number of candidates for $k=3$ and $k=4$. On the right, the average number of candidates by instance for $k=4$ who have a $100\%$, at least $90\%$, at least $80\%$, at least $50\%$, or at most $10\%$ probability of getting elected under PFR.
[…]
Figure 4: Average number of committees satisfying PSC, having a positive probability of being sampled after 50 thousand samples, and how many committees are needed to get a probability mass of at least $75\%$, $90\%$, $95\%$, and $99\%$, for $k=3$ on the left and $k=4$ on the right. The black dashed line indicates the total number of committees.
5 Experiments
We conduct some small experiments to measure how many committees our algorithm in Theorem 5 would actually randomize over. For this, we follow McCune and Graham-Squire (2024) and study their dataset of Scottish local government elections. This dataset consists of 1049 multiwinner voting instances with ranked preferences with on average 4800 voters, between 4 and 14 candidates and typically a $k$ of either $3$ or $4$. See the left side of Figure 2 for an overview over the distribution of $m$ for the different values of $k$. As the original preferences are not complete, we use the dataset of completed preferences constructed by Bardal et al. (2024). In their experimental work studying proportionality in this dataset, Bardal et al. (2024) already noted that in most instances, PSC is satisfied by most committees. In our experiments, we implement the dependent rounding scheme of Gandhi et al. (2006) and sample $50.000$ committees according to it for each instance. First, for the individual probabilities given by PFR, we see on the right side of  Figure 2 that typically on average one candidate is guaranteed to be selected. On the other hand, the number of candidates with a larger than $50\%$ probability is on average around $4$ for all values of $m$, while the average number of candidates with a very low chance (at most $10\%$) to be selected grows in the number of candidates, but even for $m=11$ does not exceed $4$, showing that usually there are several candidates with a non-trivial chance to be selected (the same graph for $k=3$ is very similar, see Figure 5 in the appendix).
Next, we measure the total number of committees with a positive probability and then the number of committees in the top $75\%$ (i.e., how many committees are needed to get a probability mass of at least $75\%$, if there were four committees sampled, each with probability $25\%$ this number would be $3$), $90\%$, $95\%$, and $99\%$ and compare these to the total number of committees satisfying PSC. We display these in Figure 4 for all values of $k$ and $m$ with more than $10$ instances. In it, we can see the effect already noticed by Bardal et al. (2024), most committees, especially for smaller values of $m$ satisfy PSC. Indeed, also most of these committees satisfying PSC were also sampled by us after fifty thousand samples. For the number of committees sampled with higher probabilities, one can for instance see that on average for $k=3$ and $m=7$ one of $7$ committees will be sampled with probability at least $75\%$ while on average with $99\%$ probability it will be one of $19$, with on average there being about $25$ out of the $35$ total committees satisfying PSC.
6 Conclusion
We studied the problem of finding multiwinner voting rules that are both candidate monotone and satisfy the proportionality axiom of proportionality for solid coalitions. First, we show that the non-resolute rule selecting all committees satisfying PSC is indeed candidate monotone. Secondly, we design a probabilistic voting rule based on Phragmén’s Ordered Rule and show that this probabilistic voting rule is both candidate monotone and can be implemented to randomize only over committees satisfying PSC.
There are several open questions and directions for future work. Most importantly, whether there is a resolute voting rule satisfying PSC and candidate monotonicity is still open. Further, it is open whether the results presented here can be extended to stronger proportionality axioms (such as rank-PJR of Brill and Peters (2023)) or to the setting with weak preferences and the axiom of generalized PSC. For both, a characterization as provided by Aziz and Lee (2022) of the respective proportionality axioms is missing as well.
7 Acknowledgments
Thank you to Dominik Peters for various discussions on this topic. This work was partially supported by the Singapore Ministry of Education under grant number MOE-T2EP20221-0001.
Appendix A Additional Graphic
[…]
Figure 5: The average number of candidates by instance for $k=3$ who have a $100\%$, at least $90\%$, at least $80\%$, at least $50\%$, or at most $10\%$ probability of getting elected under PFR.